Causwell Company began 2021 with 26,000 units of inventory on hand. The cost of each unit was $5.00. During 2021 an additional 46,000 units were purchased at a single unit cost, and 36,000 units remained on hand at the end of 2021 (36,000 units therefore were sold during 2021). Causwell uses a periodic inventory system. Cost of goods sold for 2021, applying the average cost method, is $221,400. The company is interested in determining what cost of goods sold would have been if the FIFO or LIFO methods were used.

Required: 2. Determine the cost of goods sold for 2021 using the LIFO method. (Do not round intermediate calculations.)
Answer #1
Average unit cost 6.15 =221400/36000
Cost of goods available 442800 =(26000+46000)*6.15
Cost of goods purchased 312800 =442800-(26000*5)
Unit Cost of goods purchased 6.80 =312800/46000
Cost of goods sold using the LIFO method 244800 =36000*6.80